• edited by
22,967 views
44 44 votes

$A$ and $B$ are friends. They decide to meet between 1:00 pm and 2:00 pm on a given day. There is a condition that whoever arrives first will not wait for the other for more than $15$ minutes. The probability that they will meet on that day is

  1. $1/4$
  2. $1/16$
  3. $7/16$
  4. $9/16$

9 Answers

Best answer
63 63 votes

Meeting occurs if the first person arrives between $1:00$ and $1:45$ and the second person arrives in the next $15$ minutes or if both the persons arrive between $1:45$ and $2:00.$ 

Case 1:

  • $45/60$ are favourable cases and hence probability of first person arriving between $1:00$ and $1:45$ is $3/4.$ 
  • Probability of second person arriving in the next $15$ minutes $= 15/60 = 1/4$
  • So, probability of one person arriving between $1:00$ and $1:45$ and meeting the other $= 3/4 \times 1/4 \times 2 = 3/8$ $(2$ for choosing the first arriving friend$)$ 

Case 2:

  • Both friends must arrive between $1:45$ and $2:00.$ Probability $= 1/4 \times 1/4 = 1/16.$

So, probability of a meet $= 3/8 + 1/16 = 7/16$

Correct Answer: $C$

• edited by
36 36 votes

Check this

14 14 votes
probability that one person meet on that day = 15/60 = 1/4
Prob(failing to meet) = 3/4
prob(failing to meet by both the persons) = 3/4 * 3/4 = 9/16
Prob( meet on that day by both the persons) = 1-9/16 = 7/16
6 6 votes

For such questions which are also known as probability based on areas , u can find easily using grid of 2 dimension as shown below

    1/2*F F
   1/2 * F F 1/2 * F
1/2*F F  1/2 * F  
F  1/2 * F    

The favorable area for the given problem is shown above...

So no of favourable cells  =  4  +  6 * 1/2     =    7

Therefore probability that they meet  =  No of favourable cells/ No of total cells

                                                     =   7 / 16

Hence C) should be the correct option.

Answer:
Position:
Show:

Related questions

13 13 votes
2 answers 2 answers
3.9k
3.9k views
Akash Kanase asked Feb 15, 2016
3,938 views
Raju has $14$ currency notes in his pocket consisting of only Rs. $20$ notes and Rs. $10$ notes. The total money value of the notes is Rs. $230$. The number of Rs. $10$ n...
7 7 votes
2 answers 2 answers
4.8k
4.8k views
Akash Kanase asked Feb 15, 2016
4,778 views
The data given in the following table summarizes the monthly budget of an average household.$$\begin{array}{|l|c|c|} \hline \textbf{Category} & \textbf{Amount(Rs.)} \\\hl...
16 16 votes
2 answers 2 answers
3.4k
3.4k views
Akash Kanase asked Feb 15, 2016
3,430 views
If $(1.001)$$^{1259}$= $3.52$ and $(1.001)$$^{2062}$= $7.85$, then $(1.001)$$^{3321}$=$2.23$$4.33$$11.37$$27.64$
30 30 votes
7 answers 7 answers
16.0k
16.0k views
Akash Kanase asked Feb 15, 2016
16,033 views
There are eight bags of rice looking alike, seven of which have equal weight and one is slightly heavier. The weighing balance is of unlimited capacity. Using this balanc...